rssLink RSS for all categories
 
icon_red
icon_green
icon_red
icon_red
icon_blue
icon_green
icon_green
icon_red
icon_red
icon_red
icon_orange
icon_green
icon_green
icon_green
icon_green
icon_blue
icon_green
icon_orange
icon_red
icon_green
icon_red
icon_red
icon_green
icon_red
icon_red
icon_red
icon_red
icon_orange
icon_green
 

FS#1396 — FS#5393 — th2-1-6k

Attached to Project— Network
Incident
Whole Network
CLOSED
100%
The router is online. We look for the origin of the problem.
Date:  Tuesday, 03 May 2011, 02:19AM
Reason for closing:  Done
Comment by OVH - Saturday, 30 April 2011, 14:09PM

The router restarted :
System returned to ROM by s/w reset at 11:14:23 GMT Sat Apr 30 2011
(SP by bus error at PC 0x402DF924, address 0x0)


Comment by OVH - Saturday, 30 April 2011, 14:14PM

The router is currently stbale. The trafic toward free, sfr, numericable,
google, sfinx were notably impacted during the BGP routing switch.
The DSL clients on the LNS of Paris and the chtix clients at the2 were interrupted
when the server was offline.
We are going to analyse the causes of the crash.


Comment by OVH - Saturday, 30 April 2011, 14:15PM

TLB Modification exception, CPU signal 10, PC = 0x409A2F74^M
^M
^M
-Traceback= 409A2F74 409A30DC 409A3AA4 40461024 4105EAE4 4046F49C
41046754 41046740 ^M
$0 : 00000000, AT : 44AF0000, v0 : 409A3F1C, v1 : 409A30B4^M
a0 : 63FE2490, a1 : 6621D5A0, a2 : 407B88F0, a3 : 534DEB78^M
t0 : 6621D500, t1 : 45451260, t2 : 00000001, t3 : FFFF00FF^M
t4 : 41D7E978, t5 : 000098A9, t6 : 00000000, t7 : 00000000^M
s0 : 63FE2490, s1 : 456A900C, s2 : B2204CCD, s3 : 51911E90^M
s4 : FFFFFFFF, s5 : 50026A30, s6 : 00000001, s7 : 00000000^M
t8 : 41D7C15C, t9 : 4106AC70, k0 : 5A7EB5B0, k1 : 41D84D00^M
gp : 431011F0, sp : 61B59AF8, s8 : 00000000, ra : 409A30DC^M
EPC : 409A2F74, ErrorEPC : BFFFA454, SREG : 3400FF03^M
MDLO : 00000000, MDHI : 00000000, BadVaddr : 409A30BC^M
DATA_START : 0x42DDFAF0^M
Cause 00000004 (Code 0x1): TLB Modification exception^M


Comment by OVH - Saturday, 30 April 2011, 14:18PM

After the crash, the data displayed on weathermap concerning th2
are false even though the trafic flows normally. We are going
to update the scheme.


Comment by OVH - Monday, 02 May 2011, 20:32PM

After the restarting of the router, some chtix clients
have some connectivity problems on the vlan. After
investigations we finally forced a reset of the vlans of
the concerned clients , and this seems to fix the problem.
We are waiting the feedback for the confirmation.